If an inch worm travels 1 inch every 0.3 seconds, how many hours will it take to get across a 2 km cornfield in Ohio? (1 foot = 0.3048)
Answer #1

Speed of worm = distance traveled/time taken

distance traveled = 1 inch = 0.0254 m

time taken = 0.3 sec

Speed of worm = 0.0254 m/0.3 sec = (0.0254/0.3) m/s = 0.0847 m/s

Now to get across 2 km cornfield time taken by inch worm will be:

time taken = distance/speed

distance = 2 km = 2000 m

speed of inchworm = 0.0847 m/s

So,

time taken = 2000 m/(0.0847 m/s) = (2000/0.0847) sec = 23612.75 sec

We know that 1 hr = 3600 sec, So

time taken = (23612.75 sec)*(1 hr/3600 sec) = 23612.75/3600 hr = 6.56 hrs

time taken by inchworm to get across cornfield will be 6.56 hrs

Let me know if you've any query.